# second virial coefficient; quantum corrections; generalized Lennard-Jones potential; semiclassical fluids; parabolic cylinder functions

### v1

@@ 4 Conclusions
Specializing to the gLJ potential, Eq. (4), and introducing the change of variable $r\to t=\sqrt{8\beta^{*}}(r/\sigma)^{-n}$, one can express $B^{(2)}$ in terms of the parabolic cylinder functions $\mathrm{D}_{a-2}(-\sqrt{2\beta^{*}})$, $\mathrm{D}_{a-3}(-\sqrt{2\beta^{*}})$, $\ldots$, $\mathrm{D}_{a-8}(-\sqrt{2\beta^{*}})$, with $a=\frac{d-4}{n}$. This expression can be further simplified through repeated application of Eq. (12a).
In summary, we have obtained a compact and fully explicit expression for the first quantum correction to the second virial coefficient of a $d$-dimensional gLJ fluid, expressed in terms of parabolic cylinder or generalized Hermite functions. The formulation unifies the treatment of dimensionality and stiffness, provides analytic access to the limiting behaviors, and naturally yields the quantum correction to the Boyle temperature. Beyond its intrinsic theoretical interest, the approach presented here provides a systematic framework for deriving higher-order quantum corrections of relevance in quantum and semiclassical fluid theory

### v2

@@ 5 Outlook
Specializing to the gLJ potential, Equation (4), and introducing the change of variable $r\to t=\sqrt{8\beta^{*}}(r/\sigma)^{-n}$, one can express $B^{(2)}$ in terms of the parabolic cylinder functions $\mathrm{D}_{a-2}(-\sqrt{2\beta^{*}})$, $\mathrm{D}_{a-3}(-\sqrt{2\beta^{*}})$, $\ldots$, $\mathrm{D}_{a-8}(-\sqrt{2\beta^{*}})$, with $a=\frac{d-4}{n}$. This expression can be further simplified through repeated application of Equation (12a).
The systematic nature of our approach—reducing complex integrals to combinations of parabolic cylinder functions—extends naturally to arbitrary orders. This provides a practical computational framework for exploring the convergence properties of the quantum expansion and for determining when higher-order terms become significant. Such analysis would be particularly relevant for light atoms like helium at temperatures below $\sim 50$ K, where the ratio $qB_{q}(T^{*})/4B_{c}(T^{*})$ approaches unity and second-order corrections are no longer negligible.
6 Conclusions
In this paper, we have derived an explicit and compact expression, Equation (11), for the first quantum correction to the second virial coefficient of a $d$-dimensional fluid composed of particles interacting through the gLJ $(2n,n)$ potential defined in Equation (4). As in the classical case, Equation (8), the first quantum correction has been conveniently expressed in terms of parabolic cylinder functions. For the particular case of the sLJ fluid ($d=3$, $n=6$), the expression obtained here for $B_{q}$ [see Equation (2)] is considerably more concise than the combination of Equations (9) and (10) reported previously Zhao et al. (2025); Michels (1966).
An additional asset of the present results is that they allow one to explore the combined influence of dimensionality and stiffness on the quantum correction $B_{q}$. From Equation (11), it follows that the ratio $B_{q}/n$ depends on $d$ and $n$ only through the combination (d2)/n(d-2)/n. This implies that, at a given reduced temperature $T^{*}$, the value of $B_{q}/n$ for a $d$-dimensional fluid ($d>3$) with stiffness $n$ is identical to that of a three-dimensional fluid with an effective stiffness $n_{\text{eff}}=n/(d-2)$. In contrast, for two-dimensional fluids, $B_{q}/n$ is independent of $n$ and is given by the particularly simple expression of Equation (20), which is also applicable to any $d$ in the limit $n\to\infty$.
The knowledge of $B_{q}$ has enabled us to derive the first quantum correction to the Boyle temperature [see Equation (26)]. As illustrated by Figures 1 and 2, the general trend is that the quantum corrections to both the second virial coefficient and the Boyle temperature become more significant as the potential stiffness increases and the system dimensionality decreases.
We have applied our results to the noble gases helium, neon, and argon, demonstrating that the relative quantum correction $\delta B_{2}^{*}(T)$ decreases significantly (in absolute value) from helium to argon, primarily due to the strong dependence on the quantum parameter $q\propto(m\epsilon)^{-1}$. This application illustrates the practical utility of our compact expressions for assessing quantum effects in real physical systems.
In summary, we have obtained a compact and fully explicit expression for the first quantum correction to the second virial coefficient of a $d$-dimensional gLJ fluid, expressed in terms of parabolic cylinder or generalized Hermite functions. The formulation unifies the treatment of dimensionality and stiffness, provides analytic access to the limiting behaviors, and naturally yields the quantum correction to the Boyle temperature. Beyond its intrinsic theoretical interest, the approach presented here provides a systematic framework for deriving higher-order quantum corrections (as discussed in Section 5) of relevance in quantum and semiclassical fluid theory, and its application to noble gases demonstrates its utility for understanding quantum effects in real molecular systems. The compact analytical nature of our results also makes them particularly valuable for pedagogical purposes, providing students and researchers with transparent expressions that reveal the underlying structure of quantum corrections and facilitate the development of physical intuition about quantum effects in fluids.